OpenStudy (anonymous):

Find all solutions to the equation. 7 sin2x - 14 sin x + 2 = -5

OpenStudy (anonymous):

you mean \[7\sin^2x-14\sin x+7=0\] step 1: replace \[\sin x=u\] then you get a quadratic. step 2: solve the quadratic and get two values for "u" step 3: replace "u" back with "sin x" for each case, find the values for "x"

OpenStudy (anonymous):

When i replaced it i factored it and got (u-1)(u-1).. so are you saying it should be (sin x-1)sin x-1)?

OpenStudy (anonymous):

yes.. more precisely, \[u=1\implies\sin x=1\implies x=\sin^{-1}(1)\]

OpenStudy (anonymous):

find the values of "x" now

OpenStudy (anonymous):

since the domain is provided, there are infinite values following a recursion.. write that down.